### v1

# f(R) gravity with spacetime torsion

Abstract
The present work shows the correspondence between $f(R)$ gravity and a dual scalar-tensor theory (with an antisymmetric tensor field) when the affine connection is considered to have an antisymmetric part. It turns out that the $f(R)$ action in presence of spacetime torsion can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field through derivative coupling(s).
I Introduction
Despite a huge success of Einstein’s GR starting from the observation of Mercury perihelion to black hole observation along with the detection of gravitational waves, it has various limitations and there have been many thoughts regrading the limitations of GR. From theoretical perspective, one of the main limitations of Einstein’s GR is that the connection (associated with the spacetime metric) is $assumed$ to be symmetric in its lower indices, which in turn leads to a torsion free spacetime. On other hand, the Eisntein-Hilbert action contains only the $linear~{}power$ of Ricci scalar, while the diffeomorphism symmetry demands various combinations of spacetime curvature in the gravitational action. Thus one may argue that the Einstein-Hilbert action does not obey the full diffeomorphism symmetry of the underlying theory. As a result, it turns out that the Einstein’s GR face various challenges during different evolutionary phase of the universe, like the singularity problem (also known as the Big-Bang singularity), the issue of dark matter and dark energy etc.
In order to relax the formal assumption of symmetric connection, the first attempt was made by Cartan to include the torsion in spacetime background, which with further additions was named Einstein-Cartan theory Hehl:1976kj ; Shapiro:2001rz . In particular, it was showed that the spin angular momentum of the matter field(s) can act as a source of spacetime torsion, just like the energy is the source of spacetime curvature. In this regard, there had been a lot of interest to investigate various gravitational aspects of massless Kalb-Ramond field which is a two rank antisymmetric tensor field and has spin one duff ; kalb ; ssg_KR1 ; ssg_KR2 ; ssg_KR3 ; biswarup_KR ; sumanta_KR ; Aashish:2019ykb ; Aashish:2019zsy ; Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup . Regarding the other limitation, it is well known that Einstein–Hilbert action can be generalized by adding higher order curvature terms which naturally arise from diffeomorphism property of the action. Such terms also have their origin in String theory due to quantum corrections. f(R) Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et , Gauss-Bonnet (GB) Nojiri:2005jg ; Nojiri:2005am or more generally Lanczos-Lovelock gravity are some of the candidates in higher curvature gravitational theory. In general inclusion of higher curvature terms in the action leads to the appearance of ghost from higher derivative terms resulting into Ostragradsky instability. The Gauss–Bonnet model (a special case of Lanczos–Lovelock model) is however free of this instability due to appropriate choice of various quadratic combinations of Riemann tensor, Ricci tensor and curvature scalar. In contrast to GB model f(R) gravity model however contains higher curvature terms consisting only of the scalar curvature R. Once again just as GB model, certain classes of f(R) gravity models are free from ghost-like instability. In general f(R) model can be mapped into a scalar–tensor theory at the action level by a conformal transformation of the metric Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et . The issue of instability of the original f(R) model is now reflected in the form of the kinetic and potential terms of the scalar field in the dual scalar–tensor model, where the potential will have a stable minimum and a kinetic term with proper signature. The f(R) gravity theory earned the most attention in the arena of higher curvature gravitational theories, as the f(R) theory can naturally unify the early inflationary phase of the universe with the dark energy era Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et ; Artymowski:2014gea ; Nojiri:2003ft ; Odintsov:2019mlf ; Johnson:2019vwi ; Pinto:2018rfg ; Odintsov:2019evb ; Nojiri:2019riz ; Nojiri:2019fft ; Lobo:2008sg ; Gorbunov:2010bn ; Li:2007xn ; Odintsov:2020nwm ; Odintsov:2020iui ; Appleby:2007vb ; Elizalde:2010ts ; Cognola:2007zu . Various cosmological and astrophysical aspects of higher curvature theories are explored in Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et ; Artymowski:2014gea ; Nojiri:2003ft ; Odintsov:2019mlf ; Johnson:2019vwi ; Pinto:2018rfg ; Odintsov:2019evb ; Nojiri:2019riz ; Nojiri:2019fft ; Lobo:2008sg ; Gorbunov:2010bn ; Li:2007xn ; Odintsov:2020nwm ; Odintsov:2020iui ; Appleby:2007vb ; Elizalde:2010ts ; Cognola:2007zu ; Li:2007jm ; Carter:2005fu ; Nojiri:2019dwl ; Odintsov:2022unp ; Bamba:2020qdj ; Makarenko:2016jsy ; delaCruzDombriz:2011wn ; Chakraborty:2018scm ; Kanti:2015pda ; Kanti:2015dra ; Saridakis:2017rdo ; Cognola:2006eg .
As mentioned above that without spacetime torsion, f(R) gravity action can be mapped to a minimally coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of f(R) under consideration. However one may expect that the scenario changes in presence of spacetime torsion, i.e., if the affine connection is considered to have an antisymmetric part. This is the motivation of our present work, in particular, we want to address the correspondence between the Jordan and Einstein frame f(R) gravity in the presence of spacetime torsion. Throughout the paper, a quantity with an overbar represent the same w.r.t. the symmetric Christoffel connection, and a quantity with tilde refers to the Einstein frame quantity.
II Ricci Tensor with torsion
As we know the torsion tensor is the antisymmetric part of the Christoffel connections. We know the Riemann curvature tensor formula is given by
${R^{d}}_{abc}=\partial_{b}({\Gamma^{d}}_{ac})-\partial_{c}({\Gamma^{d}}_{ab})+{\Gamma^{e}}_{ac}{\Gamma^{d}}_{be}-{\Gamma^{e}}_{ab}{\Gamma^{d}}_{ce}$ (1)
One can express the connection through metric and contorsion using metric compatibility in a unique way as
${\Gamma^{d}}_{ab}={\bar{\Gamma}^{d}}_{ab}-{K^{d}}_{ab}$
where ${\bar{\Gamma}^{d}}_{ab}$ is symmetric in the lower indices and is given by:
${\bar{\Gamma}^{d}}_{ab}=\frac{1}{2}g^{dc}\left(\partial_{a}g_{eb}+\partial_{b}g_{ea}-\partial_{e}g_{ab}\right)$ (2)
and
${K^{d}}_{ab}=\frac{1}{2}({{T^{d}}_{ab}}-{{{T_{a}}^{d}}_{b}}-{{{T_{b}}^{d}}_{a}})$ (3)
is called the contorsion tensor. The indices are raised and lowered by means of the metric. Here it may be mentioned that the contorsion is antisymmetric in the first two indices, i.e. $K_{abc}=-K_{bac}$, while the torsion tensor $T^{a}_{bc}$ itself is antisymmetric in the last two indices. Due to Eq. (3), the Reimann tensor is written as,
$\begin{split}{R^{a}}_{bcd}&={\bar{R}^{a}}_{bcd}-\partial_{c}{K^{a}}_{bd}+\partial_{d}{K^{a}}_{bc}-{\bar{\Gamma}^{a}}_{ec}{K^{e}}_{bd}-{\bar{\Gamma}^{e}}_{bd}{K^{a}}_{ec}\\ &+{\bar{\Gamma}^{a}}_{ed}{K^{e}}_{bc}+{\bar{\Gamma}^{e}}_{bc}{K^{a}}_{ed}+{K^{a}}_{ec}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{bc}\end{split}$
$\begin{split}R_{bd}&=\bar{R}_{bd}-\partial_{a}{K^{a}}_{bd}+\partial_{d}{K^{a}}_{ba}-{\bar{\Gamma}^{a}}_{ea}{K^{e}}_{bd}-{\bar{\Gamma}^{e}}_{bd}{K^{a}}_{ea}\\ &+{\bar{\Gamma}^{a}}_{ed}{K^{e}}_{ba}+{\bar{\Gamma}^{e}}_{ba}{K^{a}}_{ed}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}\end{split}$
This can also be written as
$R_{bd}=\bar{R}_{bd}-\bar{\nabla}_{a}{K^{a}}_{bd}+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}~{}~{}.$ (4)
As the covariant derivative can be expanded in the terms of the christoffel connections and its derivatives and then the two terms cancel out giving back the expression found earlier for the ricci tensor $R_{ab}$.Hence, equation (2) gives the expression for the Ricci tensor as a function of the symmetric christoffel connections and the contorsion tensor. This turns out to be sum of the Ricci tensor builded from the symmetric part of the christoffel connections and derivative of contorsion and higher order terms in it. We shall use this expression in the next section to modify the f(R) action in terms of torsion and metric coefficients.
III f(R) action with torsion in Einstein frame
As mentioned in the introduction, we intend to map the f(R) action from Jordan to Einstein frame in presence of spacetime torsion. The f(R) action can be written as:
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}f(R)~{}~{}.$ (5)
Here $f(R)$ is an analytic function of Ricci scalar: $R=R_{\mu\nu}g^{\mu\nu}$ where Ricci tensor $R_{\mu\nu}$ contains the torsional part and given by Eq.(4). Moreover $g$ is the determinant of the metric $g_{\mu\nu}$ and $\kappa^{2}=8\pi G$ with $G$ being the Newton’s gravitational constant. Owing to Eq.(4), the action (5) can be shown as,
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}~{}f\left[g^{bd}(\bar{R}_{bd}-\bar{\nabla}_{a}{K^{a}}_{bd}+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\right]~{}~{}.$ (6)
Moreover, by introducing an auxiliary field $A(x)$, the action (5) can be equivalently written as,
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}\left[f^{\prime}(A)\left(R-A\right)+f(A)\right]~{}~{}.$ (7)
The variation of this action over the auxiliary field $A(x)$ leads to $A=R$ which finally results to the original action (5). The above action can be mapped to the Einstein frame by applying the following conformal transformation on the metric $g_{\mu\nu}$:
$g_{\mu\nu}\longrightarrow\widetilde{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}~{}~{},$ (8)
where $\Omega(x)$ is the conformal factor which is related to the auxiliary field as $\Omega^{2}=f^{\prime}(A)$, and from onwards, the quantities with an over-tilde represent that in the Einstein frame. Owing to this transformation the symmetric connection and the corresponding Ricci scalar (i.e without torsion) transform as:
$\bar{\Gamma}^{\mu}_{\nu\sigma}=\widetilde{\bar{\Gamma}}^{\mu}_{\nu\sigma}+\delta_{\nu}^{\mu}\partial_{\sigma}w+\delta_{\sigma}^{\mu}\partial_{\nu}w-\tilde{g}_{\nu\sigma}\partial^{\mu}w$ (9)
and
$\bar{R}=\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)$ (10)
respectively, (recall that the symmetric connection and the corresponding quantities without torsion are denoted by an overbar). Here $w\equiv ln\Omega$ and $\widetilde{\Box}$ is the d’Alembertian operator formed by $\widetilde{g}_{\mu\nu}$, in particular
~$\widetilde{\Box}w\equiv\frac{1}{\sqrt{-\widetilde{g}}}\partial_{\mu}(\sqrt{-\widetilde{g}}~{}\widetilde{g}^{\mu\nu}\partial_{\nu}w)~{}~{}.$
Eq.(4), with the help of Eq.(9) and Eq.(10), leads to the transformation for Ricci scalar in presence of torsion. If $R$ and $\widetilde{R}$ symbolize the Ricci scalars in the Jordan and in the Einstein frame respectively, then they are related by,
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\widetilde{g}^{bd}\Omega^{2}[-\partial_{a}{K^{a}}_{bd}-(\widetilde{\Gamma}^{a}_{ea}+\delta_{a}^{e}\partial_{a}w+\delta_{a}^{a}\partial_{e}w-\widetilde{g}_{ea}\partial^{a}w)K^{e}_{bd}\\ &-(\widetilde{\Gamma}^{e}_{bd}+\delta_{b}^{e}\partial_{d}w+\delta_{d}^{e}\partial_{b}w-\widetilde{g}_{bd}\partial^{e}w){K^{a}}_{ea}+\partial_{d}{K^{a}}_{ba}+(\widetilde{\Gamma}^{a}_{ed}+\delta_{e}^{a}\partial_{d}w+\delta_{d}^{a}\partial_{e}w-\widetilde{g}_{ed}\partial^{a}w){K^{e}}_{ba}\\ &+(\widetilde{\Gamma}^{e}_{ba}+\delta_{b}^{e}\partial_{a}w+\delta_{a}^{e}\partial_{b}w-\widetilde{g}_{ba}\partial^{e}w){K^{a}}_{ed}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}]~{}~{}.\end{split}$ (11)
Actually the first term of Eq.(4) transforms as per Eq.(10), and for other terms, we expand the covariant derivatives in terms of the Christoffel connection and then use the expression of Eq.(9). On simplifying Eq.(11), we get :
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}+\widetilde{g}^{bd}(-\partial_{a}{K^{a}}_{bd}+\partial_{d}{K^{a}}_{ba}\\ &+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}-{\widetilde{\Gamma}^{a}}_{ea}{K^{e}}_{bd}-{\widetilde{\Gamma}^{e}}_{bd}{K^{a}}_{ea}+{\widetilde{\Gamma}^{a}}_{ed}{K^{e}}_{ba}+{\widetilde{\Gamma}^{e}}_{ba}{K^{a}}_{ed})\big{]}~{}~{}.\end{split}$ (12)
In the above equation after adding and subtracting ${\widetilde{\Gamma}^{e}}_{ad}{K^{a}}_{be}$ it can be written as :
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)\\ &+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}+\widetilde{g}^{bd}(-\bar{\nabla}_{a}{K^{a}}_{bd}+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\big{]}~{}~{}.\end{split}$ (13)
Now using antisymmetry of K in the first two index and the metric compatibility property  [^1] The breaking of the metric compatibility means that one needs to add one more tensor to the affine connection. This term is called non-metricity. However, we will not consider the theories with non-metricity here. the above equation boils down to:
$\bar{R}=\Omega^{2}\left[\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}+2\bar{\nabla}_{d}{K^{ad}}_{a}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\right]~{}~{}.$ (14)
Now the 5th term in the above is a total divergent term and by using Gauss divergence theorem we can say that it vanishes at the boundary. Hence the final expression for the relation between $\bar{R}$ and $\widetilde{\bar{R}}$ is:
$\bar{R}=\Omega^{2}\left[\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\right]~{}~{}.$ (15)
Owing to the above expression of $\bar{R}$, along with the relation $\sqrt{-g}=\Omega^{-4}\sqrt{-\widetilde{g}}$, the action (7) turns out to be,
[…]
$+$ $\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}-\Omega^{-4}\left(Af^{\prime}(A)-f(A)\right)\bigg{]}~{}~{}.$
The relation between $\Omega(x)$ and $A(x)$, i.e $\Omega^{2}=f^{\prime}(A)$, makes the coefficient of $\widetilde{\bar{R}}$ unity and thus the action in the Einstein frame is given by
[…]
$+$ $\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}-\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{},$
where the suffix ’E’ stands for the Einstein frame, and being a surface term, ~$\widetilde{\Box}w$ vanishes in the Einstein frame action. Eq. (24) can be written in terms of all lowered indices as :
[…]
$+$ $\Omega^{-2}K_{cea}K_{bfd}g^{ca}g^{be}g^{fd}-\Omega^{-2}K_{aed}K_{bfc}g^{ca}g^{be}g^{fd})-\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{}$
Due to a non-dynamical field, one can vary the above action with respect to $K_{abc}$ and get an algebraic equation in $K_{abc}$.
[…]
-- $\Omega^{-2}(\delta K_{aed}K_{bfc}+K_{aed}\delta K_{bfc})g^{ca}g^{be}g^{fd})\big{)}\bigg{]}~{}~{},$
[…]
-- $\Omega^{-2}K_{bfc}g^{cj}g^{bk}g^{fl}-\Omega^{-2}K_{aed}g^{al}g^{je}g^{kd}\big{)}\bigg{]}~{}~{},$
hence the equation of motion becomes:
$2\partial^{k}wg^{jl}=-\Omega^{-2}{K^{kd}}_{d}g^{jl}-\Omega^{-2}{K^{aj}}_{a}g^{kl}+K^{klj}+K^{ljk}$ (21)
Multiplying both sides with $g_{jl}$ we get:
${K^{a}}_{ea}=4\partial_{e}w~{}.$ (22)
Eq. (22) immediately leads to the solution of $K_{abc}$ as follows,
$K_{abc}=\left(\frac{4}{3\Omega^{2}}\right)\left\{\widetilde{g}_{ac}\partial_{b}w-\widetilde{g}_{bc}\partial_{a}w\right\}+\partial_{[b}B_{ac]}$ (23)
where $B_{[ac]}$ is a two rank antisymmetric tensor field, such that $B_{[ac]}g^{ac}=0$, and note that the right hand side of Eq. (23) respects the antisymmetric nature of $K_{abc}=-K_{bac}$. Due to such antisymmetric property of $B_{[ac]}$, the solution in Eq. (22) can be easily obtained from the other one. Eq. (23) argues that the d.o.f of the contorsion tensor is encapsulated within a scalar field ($w$) and a 2-rank antisymmetric tensor field ($B_{[ac]}$). Using the above solution of $K_{abc}$ from Eq. (23) to Eq. (18) along with a little bit of simplification yields the final form of the action in the Einstein frame as :
[…]
where we use $B_{[ac]}g^{ac}=0$. It may be noted that neither $w$ nor $B_{[ac]}$ in the action (24) are canonical, therefore, in order to make their kinetic terms canonical, let us redefine :
$w\rightarrow\phi\equiv\frac{2ln\Omega}{\kappa}$ (25)
and
$B_{[ab]}\rightarrow Z_{[ab]}\equiv\frac{\Omega^{2}B_{[ab]}}{\kappa}$ (26)
In terms of canonical fields, the action in Eq. [24] takes the following form:
[…]
-- $\frac{\kappa}{2}\widetilde{g}^{ac}\widetilde{g}^{be}\widetilde{g}^{fd}\left(\partial_{[e}\phi~{}Z_{ad]}\partial_{[f}Z_{bc]}+\partial_{[f}\phi~{}Z_{bc]}\partial_{[e}Z_{ad]}\right)$
$+$ $\frac{\kappa^{2}}{2}\widetilde{g}^{ac}\widetilde{g}^{be}\widetilde{g}^{fd}\partial_{[e}\phi~{}Z_{ad]}\partial_{[f}\phi~{}Z_{bc]}\bigg{]}~{},$
where the lower and upper indices are with respect to $\widetilde{g}_{ab}$. The above action resembles with a scalar-tensor action along with a 2-rank massless antisymmetric tensor field ($Z_{[ab]}$), where the scalar field gets non-minimally coupled with the 2-rank tensor field. Note that $\phi(x)$ acts as a scalar field with the potential $V(A(\phi))=\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}$, and the antisymmtric tensor field carries the signature of the spacetime torsion. Therefore in the context of $f(R)$ gravity in presence of spacetime torsion, the higher curvature d.o.f manifests itself as a scalar field d.o.f which couples with the massless antisymmetric tensor field (or equivalently, with the torsion field) through derivative kind of coupling. It is important to note that the scalar field and the torsion do not propagate independently, actually the torsion field gets a source term from $\phi(x)$. Clearly such coupling between the scalar and the torsion fields introduces new three or four point vertices which may have interesting phenomenological implications both in cosmology as well as in particle physics. Here it is important to note that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the terms in the second line of action (III), i.e., the 3-point interaction between $\phi$ and $Z_{[ab]}$. Being a derivative coupling, it should have significant effects during the early universe cosmological phenomena where the energy scale of the universe is of $\sim 10^{-3}$ order of the Planck scale. Here we should mention that an equivalent scalar-tensor representation of Cartan f(R) theory has been demonstrated in Inagaki:2022blm where the authors showed that, similar to normal f(R) gravity, Cartan f(R) theory can also be represented to a minimally coupled scalar-tensor theory (without any antisymmetric tensor field). Actually the formalism of the current work is different than Inagaki:2022blm , in particular, we transform the “f(R) gravity with torsion” from Jordan to Einstein frame based on a conformal transformation of metric, unlike to Inagaki:2022blm where the authors rewrite the Cartan f(R) theory in Jordan frame itself by a different fashion without considering any conformal transformation of the metric. This makes our present scenario essentially different from the earlier ones.
The massless rank-2 antisymmetric tensor field may be identified with the Kalb-Ramond (KR) field. The minimally coupled scalar-tensor theory along with a KR field (i.e., where there is no coupling between the scalar field and the KR field) proves to be useful for explaining various cosmic phenomena during early universe, as showed by some of our authors in Elizalde:2018rmz ; Paul:2022mup . In particular, the Kalb-Ramond is able to trigger a bouncing universe or an inflationary universe depending on the initial conditions (see Elizalde:2018rmz ; Paul:2022mup ). Regarding the inflation, it turns out that the KR field energy density decays at a faster rate compared to the radiation and the pressureless dust with the expansion of the universe — which indicates that it has significant effects during the early universe, in particular, it enhances the tensor perturbation amplitude and hence the tensor-to-scalar ratio. Consequently the KR field gives well contributions on enhancing primordial gravitational waves generated from Bunch-Davies vacuum during the early universe. Although the KR field has significant effects during the early universe, at the same time, it is also important to realize that the $present$ stage of the universe carries practically no observable footprints of higher rank antisymmetric tensor fields including the Kalb-Ramond one. Thus a natural question arises from observable signatures of scalar, fermion, and vector d.o.f. in our Universe along with spin 2 symmetric tensor field in the form of gravity: why is our Universe is free of any perceptible signature of massless antisymmetric tensor modes? A possible explanation of this issue can be found from the minimally coupled scalar-tensor-KR theory Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup .
Thus as a whole, the Einstein frame action of $f(R)$ gravity with spacetime torsion is given by a non-minimally coupled scalar-tensor theory with a rank-2 massless antisymmetric tensor field, in particular, by the action (III) . The intriguing effects of the minimally coupled scalar-tensor-KR theory, as shown in Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup , points the larger importance of the action (III) where the scalar field gets coupled with the torsion field. Such possibilities are expected to study, in detail, in some future work.
IV Summary
In summary, we address the correspondence between the Jordan and the Einstein frame $f(R)$ gravity in presence of spacetime torsion. Without any torsion, it is well known that $f(R)$ gravity action can be mapped to a $minimally$ coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of $f(R)$ under consideration. However the scenario changes with spacetime torsion, in particular, if the affine connection is considered to have an antisymmetric part. It turns out that after the inclusion of torsion, the $f(R)$ action can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field. Actually the antisymmetric field carries the signature of the spacetime torsion. Therefore, interestingly, the scalar field (coming from the higher curvature d.o.f) and the torsion field do not propagate independently, actually the torsion field gets a source term from the scalar field. Furthermore, such interaction between the scalar and the torsion fields comes with a derivative coupling, and thus introduces a momentum dependent interaction vertex factor — which may have interesting phenomenological implications both in cosmology as well as in particle physics. In this regard, it may be mentioned that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the 3-point interaction between $\phi$ and $Z_{[ab]}$..
The rank-2 massless antisymmetric tensor field may be identified with Kalb-Ramond field which was studied a lot in the sector of cosmology, black holes and particle physics. Such intriguing effects of the KR field points the larger importance of $f(R)$ theory with spacetime torsion, where the torsion field shows a non-minimal coupling with the scalaron field. This is expected to study in some future work.